# On synchronized coded systems

### v1

@@ Abstract.
From this one sees that (2,9) follows from (2.6). $\Box$
Section Markov codes and SPO-codes
subsection Markov codes Let there be given a finite alphabet $\Sigma$. A Markov code We consider codes that are Markov codes in the sense of Keller [Ke] is given by a code $\mathcal{D}\subset\Sigma^{\mathbb{Z}}$ as state space and an irreducible 0 - 1 transition For the countable state Bernoulli case see [KSW]. matrix $T_{\mathcal{D}}(d,d^{\prime}),d,d^{\prime}\in\mathcal{D}$. The concatenation set $C(\mathcal{D})$ of a Markov code $\mathcal{D}$ is the set of points $x\in\Sigma^{\mathbb{Z}}$ such that there exists a sequence of indices $i_{k},k\in\mathbb{Z},$
@@
subsection SPO-codes For a finite alphabet $\Sigma$ let there be given a bifix code $\mathcal{F}\subset\Sigma^{+}$. Denote by $\mathcal{C}_{\mathcal{F}}$ the set of words $c\in\Sigma^{+}$ that have a proper prefix $f^{-}(c)\in\mathcal{F}$ and also a proper suffix $f^{+}(c)\in\mathcal{F}$. For a word $c\in\mathcal{F}$ we denote by $\mathring{c}$ the word that is obtained by removing from $c$ the suffix $f^{+}(c)$.
@@
such that (3.1) and (3.2) hold, is unique. The closure of $C(\mathcal{C})$, to be denoted by $X(\mathcal{C})$, is the subshift that is generated by $\mathcal{C}$.
@@
By construction (see (3.6)) the words $d\in\mathcal{D}_{\mathcal{C}}$ have a prefix $f^{-}(d)\in\mathcal{F}$. We have obtained an unambiguous Markov code with state space $\mathcal{D}_{\mathcal{C}}$ and transition matrix $T_{\mathcal{D}_{\mathcal{C}}}$ that is given by
@@
In imposing the hypothesis (3.5) we have followed the approach of [KSW, Theorem 3].
section Synchronization
subsection The SP0-code of a synchronized subshift
Given a synchronized subshift $X\subset\Sigma^{\mathbb{Z}}$ we set
@@
Proof.   Apply Lemma 2.1.. $\Box$
@@
Proof.   Apply Proposition 3.4. $\Box$
section A class of synchronized shifts
We give an example such that $\partial_{Markov}(X(\mathcal{C}))$ contains finitely many orbits and an example of an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ contains a countable infinity of orbits. The SPO-coded shifts of the examples satisfy Condition H, and are therefore intrinsically ergodic by Theorem 4.7.
@@
subsection Example 1
We construct an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ is a finite set of orbits. For the construction we choose a finite set $P$ of periodic points of $(\Sigma^{\mathbb{Z}},S_{\Sigma^{\mathbb{Z}}})$.
@@
$R=prod_{p\in Pperiod(p)\textperiodcentered}.$
@@
subsection Example 2
We construct an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ is countable infinity of orbits. We denote by $p^{(k)}$ the point of period $k$ in $(\{0,1\}^{\mathbb{Z}},S_{\{0,1\}^{\mathbb{Z}}})$ that is given by
@@
section A class of coded systems
@@
remark For a word $a\in\mathcal{L}(X)$ that does not have a word in $\mathcal{C}$ as a subword one can find a word $b\in\mathcal{L}(X)$ that has the word $a$ as a prefix and that has a word in $\mathcal{C}$ as a subword: We can distinguish four cases.
@@
$\alpha_{q}=\alpha^{-}_{r_{q}},\qquad(1\leq q\leq Q).\qed$
@@
Proof.   The subshift $X$ cannot have a semisynchronizing word. This follows by Lemma (2.1) from the remark by an application of Lemma 4.1 and of Lemma 4.2. $\Box$

### v2

@@ 2. Preliminaries
From this one sees that (7) follows from (4). $\Box$
3. Markov codes and SPO-codes
3.1 $\mathbb{Markov\ codes.}$ Let there be given a finite alphabet $\Sigma$. A Markov code 11 We consider codes that are Markov codes in the sense of Keller [Ke] is given by a code $\mathcal{D}\subset\Sigma^{\mathbb{Z}}$ as state space and an irreducible 0 - 1 transition2 2For the countable state Bernoulli case see [KSW]. matrix $T_{\mathcal{D}}(d,d^{\prime}),d,d^{\prime}\in\mathcal{D}$. The concatenation set $C(\mathcal{D})$ of a Markov code $\mathcal{D}$ is the set of points $x\in\Sigma^{\mathbb{Z}}$ such that there exists a sequence of indices $i_{k},k\in\mathbb{Z},$
@@ 3. Markov codes and SPO-codes
3.2. $\mathbb{SPO-codes.}$ For a finite alphabet $\Sigma$ let there be given a bifix code $\mathcal{F}\subset\Sigma^{+}$. Denote by $\mathcal{C}_{\mathcal{F}}$ the set of words $c\in\Sigma^{+}$ that have a proper prefix $f^{-}(c)\in\mathcal{F}$ and also a proper suffix $f^{+}(c)\in\mathcal{F}$. For a word $c\in\mathcal{F}$ we denote by $\mathring{c}$ the word that is obtained by removing from $c$ the suffix $f^{+}(c)$.
@@
such that (8) and (9) hold, is unique. The closure of $C(\mathcal{C})$, to be denoted by $X(\mathcal{C})$, is the subshift that is generated by $\mathcal{C}$.
@@
By construction (see (12)) the words $d\in\mathcal{D}_{\mathcal{C}}$ have a prefix $f^{-}(d)\in\mathcal{F}$. We have obtained an unambiguous Markov code with state space $\mathcal{D}_{\mathcal{C}}$ and transition matrix $T_{\mathcal{D}_{\mathcal{C}}}$ that is given by
@@
In imposing the hypothesis (11) we have followed the approach of [KSW, Theorem 3].
4. Synchronization
4.1. $\mathbb{The\ SP0-code\ of\ a\ synchronized\ subshift}.$ Given a synchronized subshift $X\subset\Sigma^{\mathbb{Z}}$ we set
@@ 4. Synchronization
Proof.   Apply Lemma 1. $\Box$
@@
Proof.   Apply Proposition 3. $\Box$
5. A class of synchronized shifts
We give an example such that $\partial_{Markov}(X(\mathcal{C}))$ contains finitely many orbits and an example of an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ contains a countable infinity of orbits. The SPO-coded shifts of the examples satisfy Condition H, and are therefore intrinsically ergodic by Theorem 8.
@@ 5. A class of synchronized shifts
5.1. $\mathbb{Example\ 1.}$ We construct an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ is a finite set of orbits. For the construction we choose a finite set $P$ of periodic points of $(\Sigma^{\mathbb{Z}},S_{\Sigma^{\mathbb{Z}}})$.
@@
$R=\prod_{p\in P}\text{period of}(p).$
@@
5.2. $\mathbb{Example\ 2.}$ We construct an SPO-code $\mathcal{C}$ such that $\partial_{Markov}(X(\mathcal{C}))$ is countable infinity of orbits. We denote by $p^{(k)}$ the point of period $k$ in $(\{0,1\}^{\mathbb{Z}},S_{\{0,1\}^{\mathbb{Z}}})$ that is given by
@@
6. A class of coded systems
@@
Remark: For a word $a\in\mathcal{L}(X)$ that does not have a word in $\mathcal{C}$ as a subword one can find a word $b\in\mathcal{L}(X)$ that has the word $a$ as a prefix and that has a word in $\mathcal{C}$ as a subword: We can distinguish four cases.
@@
$\alpha_{q}=\alpha^{-}_{r_{q}},\qquad(1\leq q\leq Q).$
@@
Proof.   The subshift $X$ cannot have a semisynchronizing word. This follows by Lemma 1 from the remark by an application of Lemma 9 and of Lemma 10. $\Box$